OpenStudy (sodapop):

1. cos θ = 2. sin θ =

OpenStudy (sodapop):

Please Help

OpenStudy (anonymous):

Do you have more data?

OpenStudy (sodapop):

Nope

OpenStudy (anonymous):

what subject is it? The only answer i could give with such little information would be an identity. \[\sin \theta=\frac{ 1 }{ \csc \theta}\] \[\cos \theta=\frac{ 1 }{ \sec \theta }\] Thos are reciprocal identities

OpenStudy (anonymous):

Those*

OpenStudy (sodapop):

This is Algebra 2 and thank you.